OpenStudy (anonymous):

PLEASE HELP ME Solve y = x^2 + 5x + 6 by completing the square?

OpenStudy (cwrw238):

x^2 + 5x = -6 (x + 2.5)^2 - 2.5^2 = -6 (x + 2.5)^2 = -6 + 6.25 = 0.25 can you continue?

OpenStudy (cwrw238):

you can also solve this by factoring

OpenStudy (anonymous):

@cwrw238 I see what you did there but now im a little lost. how do you solve it by factoring?

OpenStudy (anonymous):

@TheSmartOne can you explain it to me?

TheSmartOne (thesmartone):

\[(x+2.5)^2=.25\] Square root both sides \[\sqrt{(x+2.5)^2}=\sqrt{.25}\]

TheSmartOne (thesmartone):

\[(x+2.5)=\pm.5\] Subtract 2.5 on both sides... \[x=-2.5\pm.5\]

OpenStudy (anonymous):

@TheSmartOne thanks!

TheSmartOne (thesmartone):

No problem...